# Evaluating Impacts of Traffic Regulations in Complex Mobility Systems Using Scenario-Based Simulations

### v2

@@ 5.3 Result assessment
In this section, we present the outcomes of the simulation scenarios under different access pricing policies and behavioural settings. To explore the potential effects of alternative parameter configurations, we construct a set of scenarios and systematically compared their impacts. The purpose of this analysis is not primarily to evaluate specific policies, but rather to illustrate a methodological approach for measuring variations in the model outputs under different conditions with the proposed methodology. The focus is on demonstrating how scenario-based simulations can be used to quantify and analyze changes in key performance indicators.
@@ 5.3.1 Scenario definition
Moreover, we incorporate scenarios where the policy targets different users based on their socio-economic characteristics, or given the differences in their vehicle emission levels. The goal of the comparison is to assess the ability of the policy to account for the differences among vehicle or user characteristics (e.g., socio-economic group, vehicle class) and to what extent the relevant indices are impacted by this group-differentiated policy. Focusing on emission reductions, we consider variations in which the price is differentiated by vehicle emission class, with a ticket fee of €10 for Euro-class levels 0–3 and a reduced fee of €5 for the remaining levels 4–6 (scenario A5). Exploring user exemptions, we differentiate per socio-economic level of the population and exempted users with economic fragility (i.e., yearly income <<13k euros), corresponding to 10% of the population (scenario A6).
To assess the sensitivity of system outcomes to behavioral assumptions, we compare scenarios in which the same daily fixed-price policy is applied to 100% of the users while adjusting key behavioral parameters to capture extreme cases. These include a highly-reduced cost acceptability, where the ticket fee is fixed at €5 but the acceptable price is reduced to be <<€0.5 (scenario B1). Then, we consider the removal of time-adjustment flexibility (scenario B2) and the removal of modal shifts (scenario B3), which prevent considering time- and mode- variability, respectively, as acceptable behavioral strategies. These three extreme cases each strongly disincentivize a specific behavioral strategy, thereby allowing us to isolate and analyze the effect of excluding that strategy. This approach is particularly informative because behavioral strategies are difficult to model and their associated parameters are often subject to substantial uncertainty. By examining the extreme case in which a given strategy is absent, we effectively explore a boundary state of the system.
@@
Table 1: Parameter configuration of the scenarios A1–A6, assessing changes in policy impacts given changes in control parameters. The remaining model parameters are set equal for all the scenarios: among behavioral parameters, $C_{50\%}=4-7$€, $\Delta_{a50\%}=\Delta_{p50\%}=1$ hours, $\Delta_{ar50\%}=\Delta_{pr50\%}=1.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$
ID Description $C_{50\%}$ [€] $\Delta_{a50\%}[h]$ $\Delta_{p50\%}[h]$ Shift to PT
@@
Table 2: Parameter configuration of the scenarios B1–B3, assessing changes in policy impacts given changes in behavioral parameters. The remaining model parameters are set equal for all the scenarios: among control parameters, $t_{s}=$ 8:$0$ AM, $t_{e}=$ 6:00 PM, $C_{0}=5$€, $F_{e}=0\%$; among other behavioral parameters, $\Delta_{ar50\%}=1.5$ hours, $\Delta_{pr50\%}=1.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$
@@ 5.3.2 Evaluation setup
The comparison between scenarios is assessed by qualitative and quantitative analysis. First, we concentrate on a set of global metrics that summarize overall system behavior. They include the total number of vehicles entering the area per day (i.e., daily inflow), the maximum number of vehicles circulating in the area during the day (i.e., max daily traffic), and the maximum number of vehicles circulating in the area during the policy activation hours (i.e., max traffic in policy hours). Note that these quantities differ in the measurement units: while inflows measure the cumulative number of entering vehicles during the whole day, the maximum traffic accounts for the maximum number of vehicles circulating at the same time in the interest period. Moreover, the analysis considers the total grams of NOx emitted by the vehicles circulating in the area during the day (i.e., daily emissions). Other than these key mobility indicators, other metrics are reported to better describe the what-if scenarios, such as the absolute number of vehicles changing their habits with a time-shift, with a modal-shift, or in another way (i.e., lost). Finally, the total earnings collected with the access tickets are summarized in the global indicator of daily revenue.
@@ 5.3.3 Findings
For the proposed scenarios A1–A6, Table 3 collects the measured value of the global indicators. As expected, the results demonstrate that when access policies are activated, there is a significant reduction in inbound flows, traffic, and emissions. It is particularly interesting to quantify how the intensity of these reductions varies depending on the applied policy. Scenarios A1–A4 highlight how the policy duration significantly affects traffic and emissions. Scenario A2 results in a larger reduction in daily inflow with respect to A1 (-56.5k vs -44.6k) and in traffic during policy hours (-5.0k vs -1.5k), with a more pronounced decrease in emissions (-209.8k NOx vs -166.7k NOx), but also a substantial reduction in maximum traffic during the day (-1.5k) in contrast to the increase in the A1 scenario (+1.4k). Scenarios A3 and A4 illustrate the effect of timing: the morning policy in A3 produces smaller impacts on inflow (-17.0k) and traffic (-0.6k policy-hours), whereas the afternoon policy in A4 modifies daily maximum traffic moderately (+0.9k) and maximum traffic in policy hours significantly (-4.4k), with emissions (-10.1% vs -11.9%) and revenue (112k in both scenarios) effects more balanced.
@@
Focusing on the changes in the behavioural assumptions, the global indicators relative to scenarios B1–B3 are reported in Table 4. All scenarios see a reduction in daily inflows and NOx emissions, but the impact on peak traffic and user behavior differs significantly. Scenario B1 shows the largest reductions in inflows (-43%) and emissions (-44%), but peak traffic slightly increases due to time-shifting, suggesting that overall demand reduction is not sufficient to prevent congestion during critical hours. The very low revenues (>>€700/day) indicate minimal financial uptake, despite substantial adjustments by a subset of users. Scenario B2 achieves a moderate decrease in inflows (-32%) and emissions (-32%), with a clear reduction in peak traffic (-19% in policy hours). This indicates that rigidity and mode-shift alone can effectively mitigate congestion while maintaining significant revenue (€252.5k/day). Finally, scenario B3 shows limited environmental gains (-15% in emissions) and increased peak traffic, reflecting the concentration of demand in fewer hours. Revenues are highest (€271.8k/day), but congestion and lost trips remain substantial.
@@
Table 3: Comparison of global indicators in the scenarios A1–A6 with respect to the baseline situation
@@
Daily emissions variation [kg NOx] -273.3k (-44%) -199.9k (-32%) -91.9k (-15%)
@@
Table 4: Comparison of global indicators in the scenarios B1–B3 with respect to the baseline situation
@@
Figure 6 illustrates the distribution of the quantities over time for the scenarios B1–B3. The results reveal how behavioral constraints fundamentally reshape traffic patterns under mobility pricing. In scenario B1, traffic, inflow, and emissions become highly volatile with extreme peaks as travelers concentrate trips to avoid fees, followed by near-zero periods in policy activation hours. Scenario B2, on the other hand, produces the most stable outcomes, with sustained moderate flows and emissions, and no high peaks near policy hours. The inability to shift from car to bus (scenario B3) generates an intermediate pattern with significant peaks close to policy activation hours, but maintained baseline traffic within policy hours. The examination of these scenarios highlights the strong dependence of system-level outcomes on behavioral flexibility assumptions, supporting the importance of comprehensive sensitivity analysis. In particular, small changes in travelers’ ability to reschedule trips or switch modes can lead to qualitatively different temporal traffic profiles and emission dynamics. These findings underscore that policy evaluations based on a single behavioral specification may be misleading and that robust mobility policy assessment requires explicit consideration of heterogeneous traveler responses.
Figure 4: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1–A4
Figure 5: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1, A5–A6
Figure 6: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios B1–B3
@@ 6 Conclusions
In this study, we present a model-driven, simulation-based approach for the ex-ante evaluation of urban traffic regulation policies, capable of capturing both direct impacts, such as traffic conditions, modal shifts, and emissions, and indirect impacts, including social and economic effects related to mobility. The application of the framework to the real-world case study of Bologna demonstrates that the integration of a physical layer, representing the transportation system and traffic flows, with a social layer, capturing individual behaviors and adaptation to policy interventions, enables systematic exploration of “what-if” scenarios and assessment of the potential consequences of alternative regulatory designs before their actual implementation. The proposed approach addresses the complexity inherent in socio-technical urban mobility systems, accounting for nonlinear interactions, uncertainty, and dynamic adaptation of user behavior, while supporting transparent and reproducible generation of counterfactual scenarios. In this way, it enables a rigorous assessment of traffic policies before their implementation, offering decision-makers actionable insights into potential impacts. The evaluation results show that the proposed framework effectively supports the analysis and assessment of the impacts of various possible policies and behavioral parameters. In addition, experimenting with extreme scenarios, in particular for what concerns behavioral parameters, appears to be a powerful approach to assess the sensitivity to these parameters: this sensitivity analysis is particularly relevant for the evaluation of the potential impacts of mobility policies, given the difficulty to provide precise and accurate models for behavioral parameters.
Several directions could enhance the proposed framework. From a spatial perspective, the current model works at a single-zone level and provides aggregated insights for one urban area. A natural extension would involve a multi-scale approach that divides the selected area into finer spatial units. This would enable the evaluation of the same key indicators at a sub-zone level. Such granularity would reveal localized hotspots and spatial patterns that may be hidden by aggregation. This would support more targeted interventions and allow policymakers to design zone-specific measures.
@@
This work was supported by project “Gemello Digitale: Governo e Valorizzazione del Patrimonio Dati” (project BO1.1.2.1.a) in the scope of the National Metro Plus Program and Southern Medium Cities 2021-2027.

### v3

@@ 5.3 Result assessment
In this section, we present the outcomes of the simulation scenarios under different access pricing policies and behavioral settings. To explore the potential effects of alternative parameter configurations, we construct a set of scenarios and systematically compare their impacts. The purpose of this analysis is not primarily to evaluate specific policies, but rather to illustrate a methodological approach for measuring variations in the model outputs under different conditions with the proposed methodology. The focus is on demonstrating how scenario-based simulations can be used to quantify and analyze changes in key performance indicators.
@@ 5.3.1 Scenario definition
Moreover, we incorporate scenarios where the policy targets different users based on their socio-economic characteristics, or given the differences in their vehicle emission levels. The goal of the comparison is to assess the ability of the policy to account for the differences among vehicle or user characteristics (e.g., socio-economic group, vehicle class) and to what extent the relevant indices are impacted by this group-differentiated policy. Focusing on emission reductions, we consider variations in which the price is differentiated by vehicle emission class, with a ticket fee of €10 for Euro-class levels 0–3 and a reduced fee of €5 for the remaining levels 4–6 (scenario A5). Exploring user exemptions, we differentiate per socio-economic level of the population and exempt users with economic fragility (i.e., yearly income <<€13k), corresponding to 10% of the population (scenario A6).
To assess the sensitivity of system outcomes to behavioral assumptions, we compare scenarios in which the same daily fixed-price policy is applied to 100% of the users while adjusting key behavioral parameters to capture extreme cases. These include a highly-reduced cost acceptability, where the ticket fee is fixed at €5 but the acceptable price is reduced to be less than €0.5 (scenario B1). Then, we consider the removal of time-adjustment flexibility (scenario B2) and the removal of modal shifts (scenario B3), which prevent considering time- and mode- variability, respectively, as acceptable strategies. These three extreme cases each strongly disincentivize a specific behavioral strategy, thereby allowing us to isolate and analyze the effect of excluding that strategy. This approach is particularly informative because behavioral strategies are difficult to model and their associated parameters are often subject to substantial uncertainty. By examining the extreme case in which a given strategy is absent, we effectively explore a boundary state of the system.
@@
Table 1: Parameter configuration of the scenarios A1–A6, assessing changes in policy impacts given changes in control parameters. The remaining model parameters are set equal for all the scenarios: among behavioral parameters, $C_{50\%}=4-7$€, $\Delta_{a50\%}=\Delta_{p50\%}=1$ hours, $\Delta_{ar50\%}=\Delta_{pr50\%}=1.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$.
ID Description $C_{50\%}$ [€] $\Delta_{a50\%}[h]$ $\Delta_{p50\%}[h]$ Shift to PTS
@@
Table 2: Parameter configuration of the scenarios B1–B3, assessing changes in policy impacts given changes in behavioral parameters. The remaining model parameters are set equal for all the scenarios: among control parameters, $t_{s}=$ 8:$0$ AM, $t_{e}=$ 6:00 PM, $C_{0}=5$€, $F_{e}=0\%$; among other behavioral parameters, $\Delta_{ar50\%}=1.5$ hours, $\Delta_{pr50\%}=1.5$ hours, $\beta_{0}=-1.24$, $\beta_{1}=4.5$, $\beta_{2}=-1.45$, $\beta_{3}=-0.30$, $\beta_{4}=-0.034$; for the dwell time distribution, $\bar{\tau}=20$ minutes; for the emission conversion factor, $km\nu=2.5$.
@@ 5.3.2 Evaluation setup
The comparison between scenarios, which is reported in Subsection 5.3.3, covers both qualitative and quantitative analysis. First, we concentrate on a set of global metrics that summarize overall system behavior. They include the total number of vehicles entering the area per day (i.e., daily inflow), the maximum number of vehicles circulating in the area during the day (i.e., max daily traffic), and the maximum number of vehicles circulating in the area during the policy activation hours (i.e., max traffic in policy hours). Note that these quantities differ in the measurement units: while inflows measure the cumulative number of entering vehicles during the whole day, the maximum traffic accounts for the maximum number of vehicles circulating at the same time in the interest period. Moreover, the analysis considers the total grams of NOx emitted by the vehicles circulating in the area during the day (i.e., daily emissions). Other than these key mobility indicators, other metrics are reported to better describe the what-if scenarios, such as the absolute number of vehicles changing their habits with a time-shift, with a modal-shift, or in another way (i.e., lost). Finally, the total earnings collected with the access tickets are summarized in the global indicator of daily revenue.
@@ 5.3.3 Findings
For the proposed scenarios A1–A6, Table 3 collects the measured values of the global indicators. As expected, the results demonstrate that when access policies are activated, there is a significant reduction in inbound flows, traffic, and emissions. It is particularly interesting to quantify how the intensity of these reductions varies depending on the applied policy. Scenarios A1–A4 highlight how the policy duration significantly affects traffic and emissions. Scenario A2 results in a larger reduction in daily inflow with respect to A1 (-56.5k vs -44.6k) and in traffic during policy hours (-5.0k vs -1.5k), with a more pronounced decrease in emissions (-209.8kg NOx vs -166.7kg NOx), but also a substantial reduction in maximum traffic during the day (-1.5k) in contrast to the increase in the A1 scenario (+1.4k). Scenarios A3 and A4 illustrate the effect of timing: the morning policy in A3 produces smaller impacts on inflow (-17.0k) and traffic (-0.6k policy-hours), whereas the afternoon policy in A4 modifies daily maximum traffic moderately (+0.9k) and maximum traffic in policy hours significantly (-4.4k), with emissions (-10.1% vs -11.9%) and revenue (112k in both scenarios) effects more balanced.
@@
Focusing on the changes in the behavioral assumptions, the global indicators relative to scenarios B1–B3 are reported in Table 4. All scenarios see a reduction in daily inflows and NOx emissions, but the impact on peak traffic and user behavior differs significantly. Scenario B1 shows the largest reductions in inflows (-43%) and emissions (-44%), but peak traffic slightly increases due to time-shifting, suggesting that overall demand reduction is not sufficient to prevent congestion during critical hours. The very low revenues (>>€700/day) indicate minimal financial uptake, despite substantial adjustments by a subset of users. Scenario B2 achieves a moderate decrease in inflows (-32%) and emissions (-32%), with a clear reduction in peak traffic (-19% in policy hours). This indicates that rigidity and mode-shift alone can effectively mitigate congestion while maintaining significant revenue (€252.5k/day). Finally, scenario B3 shows limited environmental gains (-15% in emissions) and increased peak traffic, reflecting the concentration of demand in fewer hours. Revenues are highest (€271.8k/day), but congestion and lost trips remain substantial.
@@
Table 3: Comparison of global indicators in the scenarios A1–A6 with respect to the baseline situation.
@@
Daily emissions variation [g NOx] -273.3k (-44%) -199.9k (-32%) -91.9k (-15%)
@@
Table 4: Comparison of global indicators in the scenarios B1–B3 with respect to the baseline situation.
@@
Figure 6 illustrates the distribution of the quantities over time for scenarios B1–B3. The results reveal how behavioral constraints fundamentally reshape traffic patterns under mobility pricing. In scenario B1, traffic, inflow, and emissions become highly volatile with extreme peaks as travelers concentrate trips to avoid fees, followed by near-zero periods in policy activation hours. Scenario B2, on the other hand, produces the most stable outcomes, with sustained moderate flows and emissions, and no high peaks near policy hours. The inability to shift from car to bus (scenario B3) generates an intermediate pattern with significant peaks close to policy activation hours, but maintained baseline traffic within policy hours. The examination of these scenarios highlights the strong dependence of system-level outcomes on behavioral flexibility assumptions, supporting the importance of comprehensive sensitivity analysis. In particular, small changes in travelers’ ability to reschedule trips or switch modes can lead to qualitatively different temporal traffic profiles and emission dynamics. These findings underscore that policy evaluations based on a single behavioral specification may be misleading and that robust mobility policy assessment requires explicit consideration of heterogeneous traveler responses.
We remark that, for the sake of simplicity in the comparison among different what-if scenarios, the analysis of the expressiveness done so far does not cover the uncertainty in the predictions. However, all the analyses can be extended to this aspect as well as prediction uncertainty is a built-in feature of the model. Figure 7 shows the emissions for scenarios A1–A6 and B1–B3 already discussed, but enriched with uncertainty bands. Unsurprisingly, the uncertainty is higher during the hours when the policy is active. During these hours, the uncertainty depends on the predictability of the changes in the behaviors. For instance, in the case of B1, the uncertainty is low since the scenario imposes extreme sensitivity to the policy costs.
Figure 4: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1–A4.
Figure 5: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios A1, A5–A6.
Figure 6: Inflow [veh], traffic [veh], and emissions [g NOx] varying over time for the scenarios B1–B3.
Figure 7: Uncertainty bands of emission estimates [g NOx] over time for scenarios A1–A6 and B1–B3 (ordered left-to-right, top-to-bottom). Color density represents estimation confidence using a divergent scale: the central value (50), shown in red, indicates maximum confidence (the most probable value), while the color fades to white toward the boundaries (0 and 100) to signify decreasing confidence.
@@ 6 Conclusions
In this study, we present a model-driven, simulation-based approach for the ex-ante evaluation of urban traffic regulation policies, capable of capturing both direct impacts, such as traffic conditions, and indirect impacts, including environmental effects related to mobility. The application of the framework to the real-world case study of Bologna demonstrates that the integration of a physical layer, representing the transportation system and traffic flows, with a social layer, capturing individual behaviors and adaptation to policy interventions, enables systematic exploration of what-if scenarios and assessment of the potential consequences of alternative regulatory designs before their actual implementation. The proposed approach addresses the complexity inherent in socio-technical urban mobility systems, accounting for nonlinear interactions, uncertainty, and dynamic adaptation of user behavior, while supporting transparent and reproducible generation of counterfactual scenarios. In this way, it enables a rigorous assessment of traffic policies before their implementation, offering decision-makers actionable insights into potential impacts.
The evaluation results highlight the analytical strength of the proposed framework. In particular, the framework exhibits the capacity to discriminate between policy configurations that are structurally similar, showing when differences in policy design produce meaningfully different effects on the system and when they lead to similar overall outcomes. Moreover, the analysis shows that global indicators are not always sufficient to fully characterize policy effects: while in some cases aggregate measures provide an adequate assessment, in others a temporal examination is necessary to correctly interpret congestion and flow patterns. This ability to combine aggregate and temporal perspectives strengthens the robustness of the policy evaluation process. Finally, although uncertainty is not the primary focus of this study, the framework naturally accommodates it, enabling future extensions toward fully uncertainty-aware assessments. In addition, experimenting with extreme scenarios, in particular for what concerns behavioral parameters, appears to be a powerful approach to assess the sensitivity to these parameters: this sensitivity analysis is particularly relevant for the evaluation of the potential impacts of mobility policies, given the difficulty to provide precise and accurate models for behavioral parameters.
Several directions could enhance the proposed framework. From a spatial perspective, the current model works at a single-zone level and provides aggregated insights for one urban area. A natural extension would involve a multi-scale approach that divides the selected area into finer spatial units. This would enable the evaluation of the same key indicators at a sub-zone level. Such granularity would reveal localized hotspots and spatial patterns that may be hidden by aggregation, supporting more targeted interventions and allowing policymakers to design zone-specific measures.
@@
This work was supported by project “Gemello Digitale: Governo e Valorizzazione del Patrimonio Dati” (project BO1.1.2.1.a) in the scope of the National Metro Plus Program and Southern Medium Cities 2021-2027 and the “ICSC National Research Centre for High Performance Computing, Big Data and Quantum Computing” (CN00000013), under the NRRP MUR program funded by the NextGenerationEU.